Encik Ahmad's age is three times his son's age. 10 years ago, he was five times his son's age. Calculate the age of Encik Ahmad and his son respectively.
Answer by macston(5194)

.
F=father's age=3S; S=son's age
.
F-10=5(S-10)
3s-10=5S-50
40=2S
20=S ANSWER 1: the son is 20 years old.
.
F=3S=3(20)=60 ANSWER 2: Father is 60 years old.
.
CHECK:
F-10=5(S-10)
60-10=5(20-10)
50=5(10)
50=50
.